# Quantifying Observational Projection Effects with a Simulation-based hot CGM model

### v1

@@ 3.1 Mock X-ray observation
The photons are simulated in the $0.5-2.0$ keV intrinsic band with pyXsim (ZuHone & Hallman 2016), which is based on phox (Biffi et al. 2013, 2018), by assuming an input emission model where the hot X-ray emitting gas is in collisional ionization equilibrium. The spectral model computations of hot plasma use the Astrophysical Plasma Emission Code, apec [^5] APEC link https://heasarc.gsfc.nasa.gov/xanadu/xspec/manual/node134.html code (Smith et al. 2001) with atomic data from atomdb v3.0.9 (Foster et al. 2012). This model requires the plasma temperature of the gas cells (in keV), the metal abundances, the redshift $z$ and the normalization,
@@
Within pyXSim, the number of photons generated depends on the specified collecting area of the assumed X-ray telescope, its exposure time, and redshift. We generate sufficient photons by assuming a telescope with an energy-independent collecting area of $1000$ cm2 and an exposure time of $1000$ ks. The photon-list is generated in the observed frame of the X-ray emitting gas cells and is corrected to rest frame energies. The LoS direction determines the event’s position in the sky. We define the LoS along the x-axis within the lightcone. By applying a $\theta_{\rm LC\ obs}/2=5.08$ deg rotation along the z-axis and $-\phi_{\rm LC\ obs}/2=-2.32$ deg rotation along the y-axis, we centre the y-z plane at $(0,0)$ degrees. The photons generated by the gas particles are projected onto the sky; the resulting image is shown in Fig. 2.
We also project the halo and subhalo positions on the sky, as shown in Fig. 3 for a redshift slice of $z=0.3$. The X-ray events are then stacked around the projected halo/subhalo positions. The surface brightness profiles [^7] We define the surface brightness profiles in this work in units of erg/s/kpc2, consistent with Z+24, also called the luminosity profile. are calculated $\mathcal{S}(r)$ as a function of the projected radius on the sky within the $0.5-2$ keV energy band. The number of photons with rest frame energy $E$ [erg] in each radially outward bin $r$, $N(E,r)$, is weighted by the area, $\mathcal{A}$ [kpc2], of the 2D shell between $r$ and $r+\rm{d}r$, the fraction of photons collected from the source, $f_{\rm A}$ [cm2], and the exposure time, $t_{\rm exp}$ [seconds],
@@ 3.2 X-ray surface brightness profiles and their analytic modelling
where $S_{0}$ is the central surface brightness, $r_{\rm c}$ is the core radius at which the profile slope becomes steeper, and $\beta$ is the exponent quantifying the slope of the profile. In cases where the outskirt steepens more than the slope defined for the inner radii by the $\beta-$model, following Vikhlinin et al. (2006), we introduce a new slope-parameter as follows:
@@ 3.3 Prerequisite data products for quantifying projection effects
To obtain contributions from the large-scale structure, the locally correlated X-ray emission, we generate cubes and profiles with the CENsim catalog for the two cases as follows:
@@
    R200m cubes and profiles: every central galaxy within the halo catalog is assigned X-ray events within $R_{\rm 200m}$ of the parent halo. We define these profiles as the intrinsic hot gas emission profiles.
@@
    R±3Mpc, R±9Mpc and R±27Mpc cubes and profiles: all the X-ray events, intrinsic and locally extrinsic, irrespective of whether they belong to the galaxy but within $3$ Mpc, $9$ Mpc, and $27$ Mpc of the source centre, are selected in 3D. We construct the profiles and cubes to quantify the impact of contamination from the local vicinity on the intrinsic source emission.
[…]
@@ 4.1 Fitting analytic models to the intrinsic X-ray surface brightness profiles
The increasing stellar mass bins $10^{10.5-11}\ {\rm M}_{\odot}$ , $10^{11.2-11.5}\ {\rm M}_{\odot}$, and $10^{11.2-11.5}\ {\rm M}_{\odot}$ result in decreasing values of $\beta$ from $0.4$, $0.28$, and $0.25$, respectively. The same trend holds from the halo mass bins, where increasing halo mass bins: $10^{12.5-13}\ {\rm M}_{\odot}$, $10^{13-13.5}\ {\rm M}_{\odot}$, and $10^{13.5-14}\ {\rm M}_{\odot}$ have decreasing values of $\beta$ from $0.57$, $0.30$, to $0.24$, respectively. One of the reasons for the decrease is attributed to the mass-dependence of the feedback prescriptions in the TNG300 model, i.e., stellar mode dominating at lower masses, and the kinetic and thermal modes of energy injection by AGN dominating galaxies with $M_{\star}\geq 10^{10.5}\ {\rm M}_{\odot}$ (Weinberger et al. 2016). As discussed in Weinberger et al. (2018), the two AGN-related feedback channels depend on the black-hole (BH) accretion rates. For the high accretion rates, the thermal mode causes the gas cells close to the galactic centre to heat, eventually releasing the energy radiatively in X-rays. However, for lower accretion rates, the kinetic mode kicks in, depositing energy through winds and jets in random directions away from the BH. This causes gas heating and, hence, X-ray emission via cooling at a larger distance away from the galaxy centre, overall causing a flattening of the radial X-ray surface brightness profile as shown by the slope $\beta$ in Fig. 4 and 5. This reasoning is further consolidated with studies detailing the effects of the kinetic mode of BH feedback in the TNG model (Terrazas et al. 2020), its impact on the gas properties, i.e., temperature, entropy, density, and CGM fraction (Zinger et al. 2020; Davies et al. 2020), and its correlations with X-ray emission (Truong et al. 2020; Oppenheimer et al. 2020; Truong et al. 2021; Ayromlou et al. 2023). Sorini et al. (2024) show a similar trend of decreasing slope for the gas density profiles with the SIMBA (Davé et al. 2019) suite of simulations. With simulation comparison projects like camels simulations (Villaescusa-Navarro et al. 2021), it is possible to study the correlation between X-ray emission around galaxies with different feedback implementations, the exploration of which we leave for a future study.
@@
Figure 7: Effect of centrals (purple) and misclassified-centrals (green), in the stellar mass bin $M_{\star}=\big{[}10^{10.5},10^{11}\big{]}\ {\rm M}_{\odot}$, on the total X-ray surface brightness profiles (black). The total sample is constructed such that there are $1\%$ satellites and $99\%$ centrals (left panel), $10\%$ satellites and $90\%$ centrals (middle panel), and $30\%$ satellites and $70\%$ centrals (right panel). The vertical dashed line at $242$ kpc is the mean $R_{500c}$ of the $5109$ galaxies used in the mass bin with the shaded area corresponding to the minimum and maximum values. The shaded green region is the uncertainty on the misclassified central profile obtained by bootstrapping.
@@ 4.2 Locally correlated environment
In observations, one is sensitive to the complete line of sight towards the observer. Therefore, this effect can be corrected by subtracting a background emission level determined empirically from the observed surface brightness at a large distance from the halos of interest (assuming spherical symmetry in the large-scale emission). Nevertheless, given the setup used in this work, we can, for the first time, probe this effect locally around the halo and show its stark impact in stacking experiments.
Figure 10: The fraction of X-ray emission from the intrinsic hot CGM for different levels of satellite contamination (misclassified centrals) in the total galaxy sample. The fraction of X-ray emission from the intrinsic hot CGM for misclassified central contamination fractions of $0.01$ (dotted line), $0.1$ (dash-dotted line), $0.3$ (dashed line), and $0.5$ (solid line) are shown for three different mass bins: MW-like galaxies with stellar-masses in $10^{10.5-11}\ {\rm M}_{\odot}$ (left panel), $10^{11-11.2}\ {\rm M}_{\odot}$ (middle panel), and $10^{11.2-11.5}\ {\rm M}_{\odot}$ (right panel). The shaded regions around each curve correspond to the uncertainty of the mean profile obtained by bootstrapping. The vertical black solid line at $242$ kpc in the left panel, $369$ kpc in the middle panel, and $484.5$ kpc in the right panel corresponds to the mean $R_{500c}$ of the respective mass bins with the shaded area signifying the minimum and maximum values. We find that the contaminating effect of the misclassified centrals on the stacked profiles increases when the stellar mass decreases.
@@ 4.3 The Effect of Misclassified Centrals
In simulations, we completely and accurately classify central and satellite galaxies within a given stellar mass bin. Using this to our advantage, here, we make a precise prediction of the average emission arising from stacking around satellite galaxies by considering the following fraction of satellite (or misclassified centrals) contaminating the total galaxy sample: $0.01$, $0.1$ and $0.3$. We select these contamination fractions of satellites by bootstrapping over the entire satellite galaxy sample. Tab. 1 details the number of centrals and satellites in the three stellar mass bins presented here. We show our findings in Fig. 78, and 9, corresponding to stellar mass bins $10^{10.5-11}\ {\rm M}_{\odot}$, $10^{11-11.2}\ {\rm M}_{\odot}$, and $10^{11.2-11.5}\ {\rm M}_{\odot}$, respectively.
@@
Fig. 7 quantifies the effect of misclassified centrals on the total surface brightness profile, shown with the solid black line, in the MW-stellar mass bin of $10^{10.5-11}\ {\rm M}_{\odot}$. We show the integrated profile obtained by stacking only the satellite galaxies misclassified as centrals in green and the intrinsic profile due to central galaxies in purple. The point at which the profile due to misclassified centrals (green) intersects the intrinsic central galaxy (purple) profile represents the point at which the effect due to misclassified centrals contributes $\sim 50\%$ to the total (centrals and satellites) emission. For a sample with $1\%$ (left panel) and $10\%$ (central panel) contaminating satellites, the effect due to misclassified centrals dominates over the intrinsic central galaxy emission at radii $\geq 1.2\times R_{500c}$ ($\sim 300$ kpc) and $\geq 0.2\times R_{500c}$ ($\sim 84$ kpc), respectively. In a sample with $30\%$ satellites (right panel), the effect due to misclassified centrals dominates over the intrinsic central galaxy emission at radii $\geq 0.06\times R_{500c}$ ($\sim 34$ kpc).
@@
Analogously, Fig. 8 quantifies the effect of misclassified centrals on the total surface brightness profile, shown with the solid black line, for galaxies in the stellar mass bin $10^{11-11.2}\ {\rm M}_{\odot}$. For a sample with $1\%$ (left panel) contaminating satellites, we find that the effect of misclassified centrals has negligible impact within radii $\leq R_{500c}$. For $10\%$ satellite contamination (central panel), the effect due to misclassified centrals dominates over the intrinsic central galaxy emission at radii $\geq 1.6\times R_{500c}$ ($\sim 600$ kpc) and for $30\%$ satellites contamination (right panel), the effect due to misclassified centrals dominates over the intrinsic central galaxy emission at radii $\geq 0.8\times R_{500c}$ ($\sim 300$ kpc).
@@
Finally, Fig. 9 quantifies this effect for the largest mass bin we are considering, $10^{11.2-11.5}\ {\rm M}_{\odot}$. For a sample with $1\%$ (left panel), $10\%$ (centre panel) and $30\%$ satellite contamination (right panel), the effect due to misclassified centrals remains negligible at all radii $\leq R_{500c}$. More precisely, the $1\%$ satellite contamination has a negligible impact on the total surface brightness profile. The $10\%$ and $30\%$ satellite contamination dominate over the intrinsic central galaxy profile at $\geq 2\times R_{500c}$ ($\sim 1000$ kpc) and $\geq 1.4\times R_{500c}$ ($\sim 700$ kpc), respectively.
@@
Fig. 10 presents the radial fraction of the X-ray emission from the intrinsic CGM - for four different levels of misclassified centrals contamination - in three stellar mass bins: $10^{10.5-11}\ {\rm M}_{\odot}$, $10^{11-11.25}\ {\rm M}_{\odot}$, and $10^{11.25-11.5}\ {\rm M}_{\odot}$, respectively. We complement the conclusions from Fig. 7, 8, and 9 by showing that for MW-like galaxies, the effect of misclassified centrals at the lowest contamination fraction of $0.01$ results in the intrinsic central galaxy emission contributing only $\sim 62\%$ of the total emission at $R_{500c}$. This further deteriorates with the increasing fraction of misclassified centrals in the galaxy sample, where fractions of $0.1$, $0.3$, or $0.5$ result in the intrinsic emission contributing $\sim 14\%$, $\sim 5\%$ and $\sim 3\%$ of the total emission at $R_{500c}$, respectively. This effect is less pronounced for the intermediate and most massive stellar mass bins of $10^{11-11.25}\ {\rm M}_{\odot}$ and $10^{11.25-11.5}\ {\rm M}_{\odot}$ compared to the MW-like stellar mass bin. In the case of the $10^{11-11.25}\ {\rm M}_{\odot}$ mass bin, for a fraction of misclassified centrals of $0.01$, $0.1$, $0.3$, or $0.5$, the central intrinsic emission contributes $\sim 96\%$, $\sim 71\%$, $\sim 45\%$ and $\sim 33\%$ of the total emission at $R_{500c}$, respectively. Analogously, for the stellar mass bin of $10^{11.25-11.5}\ {\rm M}_{\odot}$, for a fraction of misclassified centrals of $0.01$, $0.1$, $0.3$, or $0.5$, the central intrinsic emission contributes $\sim 99\%$, $\sim 87\%$, $\sim 68\%$ and $\sim 57\%$ of the total emission at $R_{500c}$, respectively. Therefore, we present a clear trend of the increasing importance of the effect of misclassified centrals, not only due to the increasing fraction of satellite contamination in the galaxy sample but also due to the decreasing stellar mass bins.

### v2

@@ 4.1 Mock X-ray observation
The photons are simulated in the $0.5-2.0$ keV intrinsic band with pyXsim (ZuHone & Hallman 2016), which is based on phox (Biffi et al. 2013), by assuming an input emission model where the hot X-ray emitting gas is in collisional ionisation equilibrium. The spectral model computations of hot plasma use the Astrophysical Plasma Emission Code, apec [^8] APEC link https://heasarc.gsfc.nasa.gov/xanadu/xspec/manual/node134.html code (Smith et al. 2001) with atomic data from atomdb v3.0.9 (Foster et al. 2012). This model requires the plasma temperature of the gas cells (in keV), the metal abundances, the redshift $z$ and the normalisation,
@@
Within pyXSim, the number of photons generated depends on the specified collecting area of the assumed X-ray telescope, its exposure time, and redshift. We generate sufficient photons by assuming a telescope with an energy-independent collecting area of $1000$ cm2 (about 3/4 of the eROSITA field-of-view-average effective area at 1 keV) and an exposure time of $1000$ ks. The photon-list is generated in the observed frame of the X-ray emitting gas cells and is corrected to rest frame energies. The LoS direction determines the event’s position in the sky. We define the LoS along the x-axis within the lightcone. By applying a $\theta_{\rm LC\ obs}/2=5.08$ deg rotation along the z-axis and $-\phi_{\rm LC\ obs}/2=-2.32$ deg rotation along the y-axis, we centre the y-z plane at $(0,0)$ degrees. The photons generated by the gas particles are projected onto the sky; the resulting image is shown in Fig. 2.
We also project the halo and subhalo positions on the sky, as shown in Fig. 3 for a redshift slice of $z=0.3$. The X-ray events are then stacked around the projected halo/subhalo positions. The surface brightness profiles [^10] We define the surface brightness profiles in this work in units of erg/s/kpc2, consistent with Z+24, also called the luminosity profile. are calculated $\mathcal{S}(r)$ as a function of the projected radius on the sky within the $0.5-2$ keV energy band. The centre is chosen as the location of the most bound particle within the halo/subhalo as found by the FoF/subfind algorithm. The number of photons with rest frame energy $E$ [erg] in each radially outward bin $r$, $N(E,r)$, is weighted by the area, $\mathcal{A}$ [kpc2], of the 2D shell between $r$ and $r+\rm{d}r$, the fraction of photons collected from the source, $f_{\rm A}$, and the exposure time, $t_{\rm exp}$ [seconds],
@@ 4.2 X-ray surface brightness profiles and their analytic modelling
where $S_{0}$ is the central surface brightness, $r_{\rm c}$ is the core radius at which the profile slope becomes steeper, and $\beta$ is the exponent quantifying the slope of the profile. In cases where the outskirt steepens more than the slope defined for the inner radii by the $\beta-$model, following Vikhlinin et al. (2006), we introduce a new slope-parameter,
@@ 4.3 Prerequisite data products for quantifying projection effects
To obtain contributions from the large-scale structure, the locally correlated X-ray emission, we generate cubes and profiles with the CENsim catalogue for the two cases as follows:
@@
    R200m cubes and profiles: every central galaxy within the halo catalogue is assigned X-ray events within $R_{\rm 200m}$ of the parent halo. We define these profiles as the intrinsic hot gas emission profiles.
@@
    R±3Mpc, R±9Mpc and R±27Mpc cubes and profiles: all the X-ray events, intrinsic and locally extrinsic, irrespective of whether they belong to the galaxy but within $3$ Mpc, $9$ Mpc, and $27$ Mpc of the source centre, are selected in 3D spherical comoving volumes. We construct the profiles and cubes to quantify the impact of contamination from the local vicinity on the intrinsic source emission.
To quantify contributions from the emission associated with misclassified centrals, we use the SATsim catalogue in the same stellar mass bin as the central galaxies. For galaxies in the three stellar mass bins $10^{10.5-11}\ {\rm M}_{\odot}$, $10^{11-11.25}\ {\rm M}_{\odot}$, and $10^{11.25-11.5}\ {\rm M}_{\odot}$, we construct a total galaxy samples containing $N_{\rm tot}$ galaxies, of which there are $N_{\rm cen}$ centrals and $N_{\rm sat}$ satellites. The number of satellites is defined as $N_{\rm sat}=f_{\rm sat}N_{\rm tot}$, where $f_{\rm sat}$, the fraction of satellites, is varied at the values 0.01, 0.1,0.010.10.01,\ 0.1,0.01 , 0.1 , and $0.3$. The corresponding mean X-ray surface brightness profiles for a sample containing $f_{\rm sat}$ satellites is
$\bar{\mathcal{S}}_{\rm X,tot}[r]=\frac{\sum_{i=1}^{N_{\rm tot}}\mathcal{S}_{\rm X,tot}^{(i)}}{N_{\rm tot}}.$ (8)
The mean X-ray surface brightness profile components from the central galaxies, $\bar{\mathcal{S}}_{\rm X,cen}[r]$ and satellite galaxies, $\bar{\mathcal{S}}_{\rm X,sat}[r]$ contributing to $\bar{\mathcal{S}}_{\rm X,tot}[r]$ are as follows:
$\bar{\mathcal{S}}_{\rm X,cen}[r]=(1-f_{\rm sat})\times\frac{\sum_{i=1}^{N_{\rm cen}}\mathcal{S}_{\rm X,cen}^{(i)}}{N_{\rm cen}}\rm{\ and\ }\\ \bar{\mathcal{S}}_{\rm X,sat}[r]=f_{\rm sat}\times\frac{\sum_{i=1}^{N_{\rm sat}}\mathcal{S}_{\rm X,sat}^{(i)}}{N_{\rm sat}},\ \rm{respectively}.$ (9)
Therefore, $\bar{\mathcal{S}}_{\rm X,tot}[r]=(1-f_{\rm sat})\ \bar{\mathcal{S}}_{\rm X,cen}[r]+f_{\rm sat}\ \bar{\mathcal{S}}_{\rm X,sat}[r]$. We calculate the uncertainties in the total surface brightness profile of the total, central, and satellite components by bootstrapping.
@@ 5.1 Fitting analytic models to the intrinsic X-ray surface brightness profiles
The increasing stellar mass bins $10^{10.5-11}\ {\rm M}_{\odot}$ , $10^{11.25-11.5}\ {\rm M}_{\odot}$, and $10^{11.25-11.5}\ {\rm M}_{\odot}$ result in decreasing values of $\beta$ from $0.4$, $0.28$, and $0.25$, respectively. The same trend holds from the halo mass bins, where increasing halo mass bins $10^{12.5-13}\ {\rm M}_{\odot}$, $10^{13-13.5}\ {\rm M}_{\odot}$, and $10^{13.5-14}\ {\rm M}_{\odot}$ have decreasing values of $\beta$ from $0.57$, $0.30$, to $0.24$, respectively. One of the reasons for the decrease is attributed to the mass-dependence of the feedback prescriptions in the TNG300 model, i.e., stellar mode dominating at lower masses, and the kinetic and thermal modes of energy injection by AGN dominating galaxies with $M_{\star}\geq 10^{10.5}\ {\rm M}_{\odot}$ (Weinberger et al. 2016). As discussed in Weinberger et al. (2018), the two AGN-related feedback channels depend on the black-hole (BH) accretion rates. For the high accretion rates, the thermal mode causes the gas cells close to the galactic centre to heat, eventually releasing the energy radiatively in X-rays. However, for lower accretion rates, the kinetic mode kicks in, depositing energy through winds and jets in random directions away from the BH. This causes gas heating and, hence, X-ray emission via cooling at a larger distance away from the galaxy centre, overall causing a flattening of the radial X-ray surface brightness profile as shown by the slope $\beta$ in Fig. 4 and 5. This reasoning is further consolidated with studies detailing the effects of the kinetic mode of BH feedback in the TNG model (Terrazas et al. 2020), its impact on the gas properties, i.e., temperature, entropy, density, and CGM fraction (Zinger et al. 2020; Davies et al. 2020b), and its correlation with X-ray emission (Truong et al. 2020; Oppenheimer et al. 2020; Truong et al. 2021; Ayromlou et al. 2023). Sorini et al. (2024) show a similar trend of decreasing slope for the gas density profiles with the SIMBA (Davé et al. 2019) suite of simulations. With simulation comparison projects like camels simulations (Villaescusa-Navarro et al. 2021), it is possible to study the correlation between X-ray emission around galaxies with different feedback implementations, the exploration of which we leave for a future study.
@@ 5.2 Locally correlated environment
In observations, one is sensitive to the complete line of sight towards the observer. Therefore, this effect can be corrected by subtracting a background emission level determined empirically from the observed surface brightness at a large distance from the halos of interest (assuming spherical symmetry in the large-scale emission). We test the impact of a simple, conventional background subtraction, where we subtract the mean value of the surface brightness profile beyond the $R_{\rm 200m}$ (vertical purple dotted line) from the respective grey, and dashed-grey curves in Fig. 6, as is most commonly done in observations. We find that the resulting profiles after background subtraction agree well with the intrinsic profile (solid purple line) out to $R_{\rm 500c}$ (the vertical dashed-dotted line). Beyond $R_{\rm 500c}$, such a subtraction underestimates the X-ray emission at the outskirts. However, this must be further tested with larger-volume lightcones than used here (see further discussion in Sec. 6.1). Nevertheless, given the setup used in this work, we can, for the first time, probe this effect locally around the halo and show its impact in stacking experiments.
Figure 7: Effect of centrals (purple) and misclassified-centrals (green), in the stellar mass bin $M_{\star}=\big{[}10^{10.5},10^{11}\big{]}\ {\rm M}_{\odot}$, on the total X-ray surface brightness profiles (black). The total sample is constructed such that there are $1\%$ satellites and $99\%$ centrals (left panel), $10\%$ satellites and $90\%$ centrals (middle panel), and $30\%$ satellites and $70\%$ centrals (right panel). The vertical dashed line at $242$ kpc is the mean $R_{500c}$ of the $5109$ galaxies used in the mass bin with the shaded area corresponding to the minimum and maximum values. The shaded regions on the profiles are the uncertainties obtained by bootstrapping.
@@ 5.3 The effect of misclassified centrals
In simulations, we completely and accurately classify central and satellite galaxies within a given stellar mass bin. Using this to our advantage, here, we make a precise prediction of the average emission arising from stacking around satellite galaxies by considering the following fraction of satellite (or misclassified centrals) contaminating the total galaxy sample $0.01$, $0.1$ and $0.3$. We select these contamination fractions of satellites by bootstrapping over the entire satellite galaxy sample. Tab. 1 details the number of centrals and satellites in the three stellar mass bins presented here. We show our findings in Fig. 78, and 9, corresponding to stellar mass bins $10^{10.5-11}\ {\rm M}_{\odot}$, $10^{11-11.25}\ {\rm M}_{\odot}$, and $10^{11.25-11.5}\ {\rm M}_{\odot}$, respectively.
@@
Fig. 7 quantifies the effect of misclassified centrals on the total surface brightness profile, shown with the solid black line, in the MW-stellar mass bin of $10^{10.5-11}\ {\rm M}_{\odot}$. We show the integrated profile obtained by stacking only the satellite galaxies misclassified as centrals in green and the intrinsic profile due to central galaxies in purple. The point at which the profile due to misclassified centrals (green) intersects the intrinsic central galaxy (purple) profile represents the point at which the effect due to misclassified centrals contributes $\sim 50\%$ to the total (centrals and satellites) emission. For a sample with $1\%$ (left panel) and $10\%$ (central panel) contaminating satellites, the effect due to misclassified centrals dominates over the intrinsic central galaxy emission at radii $\geq 1.04\times R_{500c}$ ($\sim 252$ kpc) and $\geq 0.24\times R_{500c}$ ($\sim 59$ kpc), respectively. In a sample with $30\%$ satellites contaminants (right panel), the effect due to misclassified centrals dominates over the intrinsic central galaxy emission at radii $\geq 0.11\times R_{500c}$ ($\sim 27$ kpc).
@@
Analogously, Fig. 8 quantifies the effect of misclassified centrals on the total surface brightness profile, shown with the solid black line, for galaxies in the stellar mass bin $10^{11-11.25}\ {\rm M}_{\odot}$. For a sample with $1\%$ (left panel) contaminating satellites, we find that the effect of misclassified centrals has negligible impact within radii $\leq R_{500c}$. For $10\%$ satellite contamination (central panel), the effect due to misclassified centrals dominates over the intrinsic central galaxy emission at radii $\geq 1.1\times R_{500c}$ ($\sim 400$ kpc) and for $30\%$ satellites contamination (right panel), the effect due to misclassified centrals dominates over the intrinsic central galaxy emission at radii $\geq 0.63\times R_{500c}$ ($\sim 231$ kpc).
@@
Finally, Fig. 9 quantifies this effect for the largest mass bin we are considering, $10^{11.25-11.5}\ {\rm M}_{\odot}$. For a sample with $1\%$ (left panel), $10\%$ (centre panel) and $30\%$ satellite contamination (right panel), the effect due to misclassified centrals remains negligible at all radii $\leq R_{500c}$. More precisely, the $1\%$ satellite contamination has a negligible impact on the total surface brightness profile. The $10\%$ and $30\%$ satellite contamination dominate over the intrinsic central galaxy profile at $\geq 1.7\times R_{500c}$ ($\sim 804$ kpc) and $\geq 1.2\times R_{500c}$ ($\sim 565$ kpc), respectively.
@@
Fig. 10 presents the radial fraction of the X-ray emission from the intrinsic CGM - for four different levels of misclassified centrals contamination - in three stellar mass bins $10^{10.5-11}\ {\rm M}_{\odot}$, $10^{11-11.25}\ {\rm M}_{\odot}$, and $10^{11.25-11.5}\ {\rm M}_{\odot}$, respectively. We complement the conclusions from Fig. 7, 8, and 9 by showing that for MW-like galaxies, the effect of misclassified centrals at the lowest contamination fraction of $0.01$ results in the intrinsic central galaxy emission contributing only $\sim 51\%$ of the total emission at $R_{500c}$. This further deteriorates with the increasing fraction of misclassified centrals in the galaxy sample, where fractions of $0.1$, $0.3$, or $0.5$ result in the intrinsic emission contributing $\sim 9\%$, $\sim 3\%$ and $\sim 1\%$ of the total emission at $R_{500c}$, respectively. This effect is less pronounced for the intermediate and most massive stellar mass bins of $10^{11-11.25}\ {\rm M}_{\odot}$ and $10^{11.25-11.5}\ {\rm M}_{\odot}$ compared to the MW-like stellar mass bin. In the case of the $10^{11-11.25}\ {\rm M}_{\odot}$ mass bin, for a fraction of misclassified centrals of $0.01$, $0.1$, $0.3$, or $0.5$, the central intrinsic emission contributes $\sim 93\%$, $\sim 55\%$, $\sim 23\%$ and $\sim 11\%$ of the total emission at $R_{500c}$, respectively. Analogously, for the stellar mass bin of $10^{11.25-11.5}\ {\rm M}_{\odot}$, for a fraction of misclassified centrals of $0.01$, $0.1$, $0.3$, or $0.5$, the central intrinsic emission contributes $\sim 99\%$, $\sim 86\%$, $\sim 61\%$ and $\sim 40\%$ of the total emission at $R_{500c}$, respectively. Therefore, we present a clear trend of the increasing importance of the effect of misclassified centrals, not only due to the increasing fraction of satellite contamination in the galaxy sample but also due to the decreasing stellar mass bins.